OpenStudy (anonymous):

The length of a rectangle exceeds its width by 12 inches, and the area is 85 square inches. What are the length and width of the rectangle?

OpenStudy (anonymous):

L=17 W=5

OpenStudy (anonymous):

L=W+12 so as per condition, W(W+12)=85 W2+12w-85=0 W2+17W-5W-85=0 W(W+17)-5(W+17)=0 (W-5)(W+17)=0 W=5 OR -17 SINCE LENGTH CANNOT BE NEGATIVE SO WIDTH MUST BE POSITIVE. SO L=5+12=17